OpenStudy (anonymous):

Two gene loci, A and B, are unlinked (and thus assort independently), and alleles A and B are dominant over alleles a and b. Indicate the probabilities of producing the following. An AB gamete from an AABb individual?

OpenStudy (aaronq):

multiply the probabilities of each allele being present A is only possibility at the first locus, thus always = 1 at the second it can be B or b, so 1/2 of the time it will be B P(A)*P(B)=1*1/2=1/2